The nuclide 96Nb decays by a first-order process with a rate constant of 2.96 × 10–2 h–1. How long will it take for 86.0% of the initial amount of 96Nb to be consumed?
Answer #1

Let initial amount is A0 = 100

Then remaining amount At = 100 - 86 = 14

K * t = 2.303 * log (A0/At)

2.96*10^-2 * t = 2.303 * log (100/14)

t = 66.43 hours ...Answer